• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin Huang

    Log entry time 04:25:59 on November 17, 2008

    Entry number 249313

    This entry is a followup to: 249288

    Followups:

    keyword=Target Polarization History during Production

    There has been 4 major spin flip run session since start of production. 
    The target polarization history has been ploted below.
    
    For Nov 08 to Nov 12 (Spin Flip Session # 209), we are using only FAP 
    laser. And equilibrium pol. is stablized to 46%
    
    For Nov 12 to Nov 13 (Spin Flip Session # 222), two new comet laser has 
    been first used in lab. The equilibrium pol. jumped to 62%
    
    Since Nov 16 (Spin Flip Session # 241 & 242), target is pumped 3 Comet 
    Laser. And the correction coil is turned on to minimize AFP loss before 
    before 10PM Nov 16. However, correciton coil canncelled too much of 
    field gradiant, which seems to causing masing effect when Pol. goes 
    above 60%. After it's turned off, Pol. has been keep rising up to 65%.
    


    A copy of this log entry has been emailed to: yqiang, jpchen



    Pol. History of Session # 209



    Pol. History of Session # 222



    Pol. History of Session # 241



    Pol. History of Session # 242